Check the validity of the following two equations to express an


electromagnetic wave.


E = k Eosin(ky+ wt) ;B = i Bosin(ky + wt)


1
Expert's answer
2021-05-25T16:52:29-0400

Answer

Not possible as because of propagation of wave is given by "E\times B" (i.e y )and it is not confirmed by

Wave "=E\times B"

"=jE_oB_osin(ky+wt)sin(ky+wt)\\"

"=jE_oB_osin^2(ky+wt)"

j (cross product of k and i as k"\times" i = j) denotes the direction towards +y .

It must be in -y direction , to get desired answer.